# Complete $k$-partite entanglement measure

### v2

@@ II.3 Coarsening relation of multipartite partitions
In what follows, we denote $A|B|C|D$, $A|B|D$, $B|D$, and $A|B$ by $ABCD$, $ABD$, $BD$, and $AB$, respectively. Namely, a system without “||||” means it is a system with $1$-fineness partition.
@@
For any subsystem of $A_{1}A_{2}\cdots A_{n}$ with arbitrary partition, it can always be derived from the global system via the coarsening relations (a)-(c) or some of them. So, based on these three coarsening relations, we can analyze not only the information relation between any subsystem and the global subsystem but also the information relation between any subsystems of the global system in detail. For instance, based on these coarsening relations, we have established the complete global entanglement measure [23, 26], the complete genuine entanglement measure [25, 26], the complete multipartite quantum discord [32], the complete multipartite quantum mutual information [33] and the complete $k$-entanglement measure [27], and discuss the complete monogamy relation of these measures [23, 32, 25, 33, 26, 27], where exploring the monogamy relation of the quantum correlations is one of the fundamental tasks in the quantum resource theory [34, 35, 37, 38, 36, 39, 40, 41, 42, 43, 23, 32, 25, 44, 33, 27, 26].
III Completeness of the $k$-PEM
@@
When we deal with the various quantum correlations living in a multipartite system, the most quintessential relation in a multipartite system is indeed the the coarsening relation, i.e., the coarsening relations of type (a)-(c), since they regardless of the type of the correlations. The “completeness” of a measure for multipartite quantum correlation mainly refers to that there is a unified criterion for quantifying different subsystems or systems under different partition, which means the amount of the quantum correlations contained in different particles or particles under arbitrary partition can be compared with each other consistently and compatibly [23, 32, 25, 26, 33, 27]. It makes up for the previous bipartite measure which can only quantify the quantum correlation under the given bipartite splitting. By reviewing the key point in defining a complete measure of quantum correlation [23, 32, 25, 26, 33, 27], we can conclude that there are two steps to reveal such a completeness of a given measure: the first step is the unification condition which is mainly related to the coarsening relation of type (a), and the second one is the hierarchy condition which is defined by the coarsening relation of type (b).
@@
We now give the definitions of the unified $k$-PEM and the complete $k$-PEM based on the coarsening relation of the partitions of the system. Hereafter, $E_{(k)}(X)$ denotes $E_{(k)}(\rho^{X})$. A $k$-PEM $E_{(k)}$ is called unified if it satisfies the unification condition: (i) (symmetry) $E_{(k)}({A_{1}A_{2}\cdots A_{n}})=E_{(k)}({A_{\pi(1)}A_{\pi(2)}\cdots A_{\pi(n)}})$ for all $\rho^{A_{1}A_{2}\cdots A_{n}}\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ and any permutation $\pi$ of $\{1,2,\cdots,n\}$; (ii) (additivity) $E_{(k)}(A_{1}A_{2}\cdots A_{r}\otimes A_{r+1}A_{r+2}\cdots A_{n})=E_{(k)}(A_{1}A_{2}\cdots A_{r})+E_{(k)}(A_{r+1}A_{r+2}\cdots A_{n})$ holds for all $\rho^{A_{1}A_{2}\cdots A_{r}}\otimes\rho^{A_{r+1}A_{r+2}\cdots A_{n}}$; (iii) ($k$-monotone)
@@
holds for all $\rho^{A_{1}A_{2}\cdots A_{n}}\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $k\geqslant 3$; and (iv) (coarsening monotone)
@@
holds for all states $\rho\in\mathcal{S}^{X_{1}X_{2}\cdots X_{p}}$ whenever $X_{1}|X_{2}|\cdots|X_{p}\succ^{a}Y_{1}|Y_{2}|\cdots|Y_{q}$ with $k\leqslant q\leqslant p$. Item (i) is clear, i.e., the symmetry is an inherent feature of any entanglement measure indeed. The $k$-partite entanglement contained in $A_{1}A_{2}\cdots A_{r}\otimes A_{r+1}A_{r+2}\cdots A_{n}$ is composed of two parts, i.e., $A_{1}A_{2}\cdots A_{r}$ and $A_{r+1}A_{r+2}\cdots A_{n}$. So we demand item (ii). If a state is $(k-1)$-producible, it must be $k$-producible, but not vice versa, so we require condition (iii). For the generalized $n$-qudit GHZ state $\frac{1}{\sqrt{d}}(|00\cdots 0\rangle+|11\cdots 1\rangle+\cdots|d-1\rangle|d-1\rangle\cdots|d-1\rangle)$, Eq. (12) is always true for any $k$-PEM. In addition, $E_{(k)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{A_{k+1}}\cdots|\psi\rangle^{A_{n}})\geqslant E_{(k)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{k-1}}|\psi\rangle^{A_{k+1}}\cdots|\psi\rangle^{A_{n}})=0$ for any $|\psi\rangle^{A_{1}A_{2}\cdots A_{k}}|\psi\rangle^{A_{k+1}}\cdots|\psi\rangle^{A_{n}}$. Therefore item (iv) is straightforward from this point of view. Hereafter, if a $k$-PEM $E_{(k)}$ obeys Eq. (11) and Eq. (12), we call it is $k$-monotonic and coarsening monotonic, respectively.
@@
holds for all $k$-partite entangled state $\rho\in\mathcal{S}^{X_{1}X_{2}\cdots X_{p}}$ whenever $X_{1}|X_{2}|\cdots|X_{p}\succ^{b}Y_{1}|Y_{2}|\cdots|Y_{q}$ with $k\leqslant q\leqslant p$. If a $k$-PEM $E_{(k)}$ satisfies Eq. (13), we call it is tightly coarsening monotonic. One need note here that, for any given $k$-PEM $E_{(k)}$,
@@
for any $\rho^{ABCD}\in\mathcal{S}^{ABCD}$, and $E_{(3)}$ is coarsening monotonic refers to
@@
for any state $\rho^{ABCD}\in\mathcal{S}^{ABCD}$. $E_{(3)}$ is tightly coarsening monotonic means
@@
IV Two classes of $k$-PEMos
In this section, we give two classes of $k$-PEMos, where the first class is similar to that of the $k$-entanglement measure defined by the minimal sum of the reduced functions in Ref. [27] and the second class is based on the unified multipartite entanglement measure introduced in Ref. [23, 24, 26].
IV.1 $k$-PEMo from the minimal sum
Let $|\psi\rangle=|\psi\rangle^{A_{1}A_{2}\cdots A_{n}}$ be a pure state in $\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$ and $h$ be a non-negative concave function on $\mathcal{S}^{X}$ with some abuse of notations. For any $\gamma_{i}^{f}\in\Gamma_{k}^{f}$, we write
where $X_{1(i)}|X_{2(i)}|\cdots|X_{m(i)}$ corresponds to $\gamma_{i}^{f}$, $\rho^{X}={\rm Tr}_{\overline{X}}|\psi\rangle\langle\psi|$, and $\overline{X}$ denotes the subsystems complementary to those of $X$. The coefficient “1/2” is fixed by the unification condition when the measures defined via $\mathcal{P}_{k}^{\gamma_{i}^{f}}$ are regarded as unified $k$-PEMs. We define
$E_{(k)}(|\psi\rangle)=\min\limits_{\Gamma_{\!k-\!1}^{f}}\mathcal{P}_{\!k-\!1}^{\gamma_{i}^{f}}(|\psi\rangle),$ (16)
where the minimum is taken over all feasible $(k\!\!-\!\!1)$-fineness partitions in $\Gamma_{\!k-\!1}^{f}$. For mixed states, we define it by the convex-roof structure. In what follows, we give only the measures for pure states, for the case of mixed states they are all defined by the convex-roof extension with no further statement. Obviously, any measure that is defined in this way is convex straightforwardly. By definition, for any $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $E_{(k)}(\rho)>0$ if and only if $\rho$ is $k$-partite entangled. Hereafter, if a measure of entanglement for pure state is defined via some function of the reduced states, such as $h$ in Eq. (15), such a function is called reduced function (also see in Ref. [44, 27, 27]).
Theorem 1.
$E_{(k)}$ is a unified $k$-PEMo. $E_{(2)}$ is complete if the reduced function is subadditive.
Proof.
By definition, it is straightforward that $E_{(k)}$ is a $k$-PEMo. We show below it satisfies the unification conditions (i)-(iv). We only need to check items (iii) and (iv) since (i) and (ii) are clear. Since $\Gamma^{f}_{k\!-\!2}\subseteq\Gamma^{f}_{k\!-\!1}$, this implies (iii) is true. For any given $|\psi\rangle\in\mathcal{H}^{A_{1}A_{2}\cdots A_{n}}$, we assume with no loss of generality that
$E_{(k)}(|\psi\rangle)=\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})+h(\rho^{X_{4}})\right]$
for some $X_{1}|X_{2}|\cdots|X_{m}$. If $Y_{1}|Y_{2}|\cdots|Y_{q}\succ^{a}Y_{1}|Y_{2}|\cdots|Y_{q}=X_{1}|X_{2}|X_{3}|X_{4}|\cdots|X_{i-1}|X_{i+1}|\cdots|X_{m}$ with $i\geqslant 6$, Eq. (12) is clear. If $Y_{1}|Y_{2}|\cdots|Y_{q}\succ^{a}Y_{1}|Y_{2}|\cdots|Y_{q}=X_{1}|X_{3}|X_{4}|\cdots|X_{m}$ with $i\geqslant 3$, we assume with no loss of generality that $|\psi\rangle^{X_{1}X_{2}X_{3}X_{4}}=|\psi\rangle^{X_{1}X_{2}}|\psi\rangle^{X_{3}}|\psi\rangle^{X_{4}}$. It turns out that
$E_{(k)}(|\psi\rangle^{X_{1}X_{2}}|\psi\rangle^{X_{3}}|\psi\rangle^{X_{4}})$
$=$ $\frac{1}{2}\left[h(\rho^{X_{1}})+h(\rho^{X_{2}})+h(\rho^{X_{3}})+h(\rho^{X_{4}})\right]$
$\geqslant$ $\frac{1}{2}\left[\sum_{i}p_{i}h(\rho^{X_{1(i)}})+h(\rho^{X_{3}})+h(\rho^{X_{4}})\right]$
$\geqslant$ $E_{(k)}(\rho^{X_{1}X_{3}X_{4}})$
for any pure state ensemble $\{p_{i},\rho^{X_{1(i)}}\}$ of $\rho^{X_{1}}$ since $h$ is concave. That is, Eq. (12) holds true still.
It is clear that the completeness of $E_{(2)}$ is reduced to the subadditivity of the reduced function. This completes the proof. ∎
If $h$ is subadditive, then the minimal partition is the ones that contained in $\Gamma_{\!k-\!1}^{f}\!\!\setminus\!\Gamma_{\!k-\!2}^{f}$. $E_{(k)}$ is not complete in general if $k\geqslant 3$. For example, if
$E_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})$
$=$ $\frac{1}{2}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right]$
it follows that
$E_{(3)}(A|BC|D|E|F|G|H)$
$=$ $\frac{1}{2}\left[h(\rho^{A})+h(\rho^{BC})+h(\rho^{FG})+h(\rho^{H})\right]$
$\geqslant$ $E_{(3)}(ABCDEFGH)$
whenever $h(\rho^{A})>h(\rho^{C})$.
$C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})$
$=$ $\frac{1}{5}\left[h(\rho^{AB})+h(\rho^{C})+h(\rho^{FG})+h(\rho^{H})\right],$
then $C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})=\frac{1}{8}(h_{A}+h_{B}+h_{C}+h_{D}+h_{E}+h_{F}+h_{G}+h_{H})$. Hereafter, we denote $h(\rho^{X})$ by $h_{X}$ for simplicity. Clearly, it is not necessary that $C_{(3)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})\leq C_{(2)}(|\psi\rangle^{ABC}|\psi\rangle^{DE}|\psi\rangle^{FGH})$. So it is not $k$-monotonic. Let $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\frac{1}{4}(h_{AB}+h_{C}+h_{DE}+h_{F})=\frac{1}{4}(h_{DE}+h_{F})$. Then $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\frac{1}{3}(h_{AB}+h_{DE}+h_{F})=\frac{1}{3}(h_{DE}+h_{F})$, i.e., $C_{(3)}$ is not coarsening monotonic. In addition $C_{(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}+C_{(3)}(|\psi\rangle^{DEF})=C_{(3)}(|\psi\rangle^{DEF})=\frac{1}{2}(h_{DE}+h_{F})$, so $C_{(3)}$ is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\frac{1}{4}(h_{A}+h_{B}+h_{C}+h_{D})=\frac{1}{2}(h_{A}+h_{C})$ may be not larger than $C_{(2)}(AB|C|D)=2h_{C}/3$.
Take $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE}$, then
$C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DE})$
$=$ $\sqrt{\frac{2}{5}}\sqrt{h_{A}+h_{D}}$
$\neq$ $C_{q(2)}(|\psi\rangle^{AB})+C_{q(2)}(|\psi\rangle^{C}|\psi\rangle^{DE})$
$=$ $\sqrt{h_{A}}+\sqrt{2h_{D}/3},$
in general. So it is not additive. For $|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF}$, we assume that $C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{DE}+h_{F}}/2$. But $C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{(h_{A}+h_{B}+h_{D}+h_{E}+h_{F})/6}$, which can not guarantee $C_{q(3)}\leq C_{q(2)}$. $C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{DEF})=\sqrt{h_{F}}/2<C_{q(3)}(|\psi\rangle^{AB}|\psi\rangle^{DEF})=\sqrt{2h_{F}/3}$ implies that $C_{q(3)}$ is not coarsening monotonic. For $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}$ may be not larger than $C_{q(2)}(AB|C|D)=\sqrt{2h_{C}/3}$, i.e., it is not tightly coarsening monotonic.
Consider $|\psi\rangle^{AB}|\psi\rangle^{CD}$, $C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{CD})=\sqrt{(h_{A}+h_{C})/2}\neq C_{G,q(2)}(|\psi\rangle^{AB})+C_{G-q(2)}(|\psi\rangle^{CD})=\sqrt{h_{A}}+\sqrt{h_{C}}$ whenever $h_{A}\neq h_{C}$. So it is not additive. Let $|\psi_{1}\rangle=|\psi\rangle^{ABC}|\psi\rangle^{D}$ with $\rho^{A}=\rho^{B}=\rho^{C}$. It turns out that $C_{G,q(3)}(|\psi_{1}\rangle)=\sqrt[20]{2h_{A}^{10}/9}>C_{G,q(2)}(|\psi_{1}\rangle)=\sqrt{3h_{A}/4}$. Namely, it is not $k$-monotonic. Let $|\psi_{2}\rangle=|\psi\rangle^{AB}|\psi\rangle^{C}|\psi\rangle^{D}$. Then $C_{G,q(2)}(|\psi_{2}\rangle)=\sqrt{2h_{A}}/2<C_{G,q(2)}(|\psi\rangle^{AB}|\psi\rangle^{D})=C_{G,q(2)}(A|B|CD)=\sqrt{2h_{A}/3}$, i.e., it is neither coarsening monotonic nor tightly coarsening monotonic.
IV.2 $k$-PEMo from unified MEM
If $|\psi\rangle=|\psi\rangle^{AB}|\psi\rangle^{CDE}|\psi\rangle^{FGH}|\psi\rangle^{I}$ with $|\psi\rangle^{CDE}$ and $|\psi\rangle^{FGH}$ are genuinely entangled, then the 3-partite entanglement is only contained in $|\psi\rangle^{CDE}$ and $|\psi\rangle^{FGH}$. The 3-partite entanglement of $|\psi\rangle$ can be quantified as $E^{(3)}(|\psi\rangle^{CDE})+E^{(3)}(|\psi\rangle^{FGH})$ for some unified multipartite entanglement measure $E^{(k)}$ (the unified multipartite entanglement measure (MEM) was introduced in Ref. [23, 26], e.g., $E^{(n)}(|\psi\rangle^{A_{1}A_{2}\cdots A_{n}})=\frac{1}{2}\sum_{i}S(\rho^{A_{i}})$ is a unified MEM, where $S(\rho)=-{\rm Tr}(\rho\log_{2}\rho)$ is the von Neumann entropy of $\rho$). Similarly, the 2-partite entanglement should be $E^{(2)}(\psi\rangle^{AB})+E^{(3)}(|\psi\rangle^{CDE})+E^{(3)}(|\psi\rangle^{FGH})$.

### v3

@@ 2.3 Coarsening relation of multipartite partitions
Here, we denote $ABCD$, ABDABD, BDBD and $AB$ by $A|B|C|D$, $A|B|D$, $B|D$ and $A|B$, respectively.
@@
For any subsystem of $A_{1}A_{2}\cdots A_{n}$ with arbitrary partition, it can always be derived from the global system via the coarsening relations (a)-(c) or some of them. So, based on these three coarsening relations, we can analyze not only the information relation between any subsystem and the global subsystem but also the information relation between any subsystems in detail. For instance, based on these coarsening relations, we have established the complete global entanglement measure [21, 24], the complete genuine entanglement measure [23, 24], the complete multipartite quantum discord [25], the complete multipartite quantum mutual information [26] and the complete $k$-entanglement measure [2]. Furthermore, we discussed the complete monogamy relation of these measures [2, 21, 23, 24, 25, 26], where exploring the monogamy relation of the quantum correlations is one of the fundamental tasks in the quantum resource theory [21, 25, 27, 28, 29, 30, 31, 32, 33, 34].
3 Completeness of the $k$-PEM
@@
When we deal with the various quantum correlations living in a multipartite system, the most quintessential relation in a multipartite system is indeed the the coarsening relation, i.e., the coarsening relations of type (a)-(c). The “completeness” of a measure for multipartite quantum correlation mainly refers to that there is a unified criterion for quantifying different subsystems or systems under different partition, which means the amount of the quantum correlations contained in different particles or particles under arbitrary partition can be compared with each other consistently and compatibly [2, 21, 23, 25, 26]. It makes up for the previous bipartite measure which can only quantify the quantum correlation under the given bipartite splitting. By reviewing the key point in defining a complete measure of quantum correlation [2, 21, 23, 25, 26], we can conclude that there are two steps to reveal such a completeness of a given measure: the first step is the unification condition which is mainly related to the coarsening relation of type (a), and the second one is the hierarchy condition which is determined by the coarsening relation of type (b).
@@
We now give the axiomatic postulates for the unified $k$-PEM and the complete $k$-PEM based on the coarsening relation of the partitions of the system. Hereafter, $E_{(k)}(X)$ denotes $E_{(k)}(\rho^{X})$. A $k$-PEM $E_{(k)}$ is called unified if it satisfies the unification condition: (i) (symmetry) $E_{(k)}({A_{1}A_{2}\cdots A_{n}})=E_{(k)}({A_{\pi(1)}A_{\pi(2)}\cdots A_{\pi(n)}})$ for all $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ and any permutation $\pi$ of $\{1,2,\cdots,n\}$; (ii) (additivity) $E_{(k)}(A_{1}A_{2}\cdots A_{r}\otimes A_{r+1}A_{r+2}\cdots A_{n})=E_{(k)}(A_{1}A_{2}\cdots A_{r})+E_{(k)}(A_{r+1}A_{r+2}\cdots A_{n})$ holds for all $\rho^{A_{1}A_{2}\cdots A_{r}}\otimes\rho^{A_{r+1}A_{r+2}\cdots A_{n}}$; (iii) ($k$-monotone)
@@
holds for all $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$, $k\geqslant 3$; and (iv) (coarsening monotone)
@@
[…]
@@
$E_{(k)}(A_{1}A_{2}\cdots A_{n})\geqslant E_{(k)}(Y_{1}|Y_{2}|\cdots|Y_{q})$ (13)
holds for all state $\rho\in\mathcal{S}^{A_{1}A_{2}\cdots A_{n}}$ whenever $A_{1}A_{2}\cdots A_{n}\succ^{b}Y_{1}|Y_{2}|\cdots|Y_{q}$ such that, for any $i$, either $\rho^{Y_{i}}$ is pure or $\rho^{Y_{i}}$ is the reduced state of some genuinely entangled pure state (or entangled bipartite pure state), $1\leqslant i\leqslant q<n$. If a $k$-PEM $E_{(k)}$ satisfies Eq. (13), we call it is tightly coarsening monotonic. For example, take $|\psi\rangle=|\psi\rangle^{ABCD}|\psi\rangle^{EF}|\psi\rangle^{GH}|\psi\rangle^{I}$ with $|\psi\rangle^{ABCD}$ is genuinely entangled, $|\psi\rangle^{EF}$ and $|\psi\rangle^{GH}$ are entangled, then the partition $AB|CD|EF|GHI$ is such a case since $|\psi\rangle^{EF}$ and $|\psi\rangle^{GHI}$ are pure states, $\rho^{AB}$ and $\rho^{CD}$ are reduced states of the genuine pure state $|\psi\rangle^{ABCD}$. That is, at least intuitively, such a coarsening operation can not generate $k$-partite entanglement. So we hope Eq. (13) should be satisfied. In addition, we take
then
$E_{(3)}(|\psi\rangle)=E_{(3)}(|\psi\rangle^{ABC}).$
But
$E_{(3)}(A|B|C|D|EFGI|H|J)=E_{(3)}(ABC)+E_{(3)}(D|EFGI|H|J)$
which is larger than $E_{(3)}(|\psi\rangle)$ whenever $|\psi\rangle^{DE}$, $|\psi\rangle^{GH}$ and $|\psi\rangle^{IJ}$ are entangled states. Namely, Eq. (13) is violated under the general coarsening operation of type (b). We thus adjust the hierarchy condition (i.e., the tight coarsening monotone condition) as Eq. (13) in stead of the “strict” hierarchy condition in Refs. [2, 21].
One need note here that, for any given $k$-PEM $E_{(k)}$,
@@
for any $\rho\in\mathcal{S}^{ABCD}$, and $E_{(3)}$ is coarsening monotonic refers to
@@
for any state $\rho\in\mathcal{S}^{ABCD}$. If $E_{(3)}$ is tightly coarsening monotonic,
@@ 4 Two classes of $k$-PEMos
4.1 $k$-PEMo from unified multipartite entanglement measure
[…]